OpenStudy (anonymous):

a 12- foot basket hoop has a shadow 9.6 ft long. how long is the shadow of a 6 ft adult standing next to the basket ball hoop

OpenStudy (calculusxy):

we can make a proportion in here \(\large \frac{\text{Height of the object}}{\text{How long the shadow is (ft)}}\)

OpenStudy (calculusxy):

We know that the basketball hoop is 12ft long (height) and its shadow is 9.6ft We also know that the height of adult is 6ft, but we don't know how long its shadow is (x) \(\huge \frac{12}{9.6} = \frac{6}{x}\)

OpenStudy (calculusxy):

Now you need to solve for x \(\large 9.6(6) = 12(x)\)

OpenStudy (calculusxy):

what do you get @meeep ?

OpenStudy (anonymous):

@calculusxy

OpenStudy (calculusxy):

what is 9.6 x 6?

OpenStudy (anonymous):

57.6

OpenStudy (anonymous):

@calculusxy